xxprettyneshh:

For nitrogen to be a liquid, its temperature must be within 12.78 °F of –333.22 °F. The equation |x + 333.22 | = 12.78 can be used to find x, which represents the maximum and minimum temperatures at which nitrogen is a liquid.

Mercury:

Next time please post the full question. It is asking for the maximum and minimum temperatures at which nitrogen is still a liquid. 2 ways you could approach this: 1. solve the equation. if |x + 333.22 | = 12.78 that means the quantity within the absolute value sign must be equal to +12.78 or -12.78 since the absolute value "removes" the negative sign. therefore, solve x + 333.22 = 12.78 and x + 333.22 = -12.78. the first equation will give you the max temperature and the second equation will give you the minimum temperature. 2. interpret the statement "its temperature must be within 12.78 °F of –333.22 °F". that means the upper bound is 12.78 more than -333.22, and the lower bound is 12.78 less than -333.22.